OpenStudy (shamil98):

∫x^2 3^x dx u = x^2 du = 2x dx dv = 3^x v = 3^x/ ln (3) + c ∫udv = uv - ∫vdu so

OpenStudy (shamil98):

x^2(3^x/ln(3) +c) - ∫ 3^x/ ln(3) +c * x^2 ?

OpenStudy (anonymous):

your trying to find the integral right?

OpenStudy (shamil98):

integrate it by parts.

zepdrix (zepdrix):

shamil, you've got the right idea, but when you go from dv to v, we don't need a constant of integration.

OpenStudy (bahrom7893):

Yea, just add the C at the very end, and here you just need to integrate by parts 2ce.

OpenStudy (shamil98):

x^2(3^x/ln(3))- ∫ 3^x/ ln(3) * x^2 + C then?

zepdrix (zepdrix):

Woops you plugged in `u` into your second part. You wanted to plug in `du`, right?

OpenStudy (shamil98):

oh oops

OpenStudy (shamil98):

x^2(3^x/ln(3))- ∫ 3^x/ ln(3) * 2x + C

zepdrix (zepdrix):

Ok yes, looks good now :)\[\Large\bf\frac{1}{\ln3}x^2 3^x-\frac{2}{\ln3}\int\limits x3^x\;dx\]

OpenStudy (shamil98):

Thank you!

zepdrix (zepdrix):

So you brought x^2 down to x, just needs parts again I guess right? :o

OpenStudy (shamil98):

Yeah, i was told i would need to integrate it twice.. not sure about that part tho.

OpenStudy (shamil98):

i just do ∫x3^x dx now?

zepdrix (zepdrix):

yah, just make sure you `distribute` the \(\Large\bf -\dfrac{2}{ln3}\) to each part of your new setup.

OpenStudy (shamil98):

Alright. so u = x du = 1 dv = 3^x v = 3^x/ln(3) + c ?

zepdrix (zepdrix):

v=3^x/ln(3)

zepdrix (zepdrix):

crapp my game is starting :c

OpenStudy (shamil98):

hold up lol ∫udv = uv - ∫vdu -2/ln3(x3^x - ∫3^x/ln(3) + C)

zepdrix (zepdrix):

\[\large\bf\frac{1}{\ln3}x^2 3^x-\frac{2}{\ln3}\left(\color{red}{\frac{1}{\ln3}}x3^x-\frac{1}{\ln3}\int\limits 3^x\;dx\right)\]So the brackets is our second by-parts. You missed this small piece in red.

OpenStudy (shamil98):

oh dang it, but thanks anyways , has been a good learning process